A DI has $19 million in T-bills, a $11 million line of credit to borrow in the repo market, and $11 million in excess cash reserves with the Fed. The DI currently has borrowed $13 million in fed funds and $9 million from the Fed discount window to meet seasonal demands.

a. What is the DI’s total available (sources of) liquidity?
b. What is the DI’s current total uses of liquidity?
c. What is the net liquidity of the DI?
d. What conclusions can you derive from the result?

(For all requirements, enter your answers in millions.)
Answer #1

a. DI’s total available (sources of) liquidity = $19 +$11 + $11 = $41 million

b.  DI’s current total uses of liquidity = 13 + 9 = $22 million

c. net liquidity of the DI = 41 - 22= $19 million

d. A net liquidity suggets that DI can withstand unexpected withdrawals of $19 million without moving the less liquid at fire sale prices.